OpenStudy (anonymous):

2(x - 3) + 2(x - 7) = 40 solve for x

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

so what does x=?

OpenStudy (anonymous):

20+10=2.x

OpenStudy (anonymous):

the choices are A. x = 12.5 B. x = 15 C. x = 5 D. no solution E. infinitely many solutions

OpenStudy (anonymous):

i dont see how that works.

OpenStudy (anonymous):

well @davidbecerra are we doing your home work buddy

OpenStudy (anonymous):

-.- cody gtfo haha

OpenStudy (anonymous):

im trying to help u shall be baned for saying that

OpenStudy (anonymous):

2x-6+2x-14=40 4x-20=40 4x=40-20 4x=20 x=50/4=5 so,x=5

OpenStudy (anonymous):

thanks @Parvathysubhash

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

it's wrong

OpenStudy (anonymous):

hi @davidbecerra